OpenStudy (anonymous):

What are the solutions of 6x2 = 12x? x = 0, x = −2 x = 0, x = 2 x = 6, x = 2 x = 6, x = −2

OpenStudy (anonymous):

Help ...!?

Parth (parthkohli):

Start by dividing both sides by 6.\[x^2 = 2x \]Now subtract 2x from both sides.\[x^2 - 2x = 0 \]Now factor, and use the zero-product rule\[x(x - 2) = 0 \]